Traffic signs are regulated by the Manual on Uniform Traffic Control Devices (MUTCD). The premiter of some traffic sign msut be 116 inches. Also, its lenght is 10 inches longer than its width. Find the dimensions of this sign. Please help me solve to obtain the width of the sign is ____ inches and the lenght of the sign is _____ inches.
Answer by stanbon(75887)

Traffic signs are regulated by the Manual on Uniform Traffic Control Devices (MUTCD). The perimeeter of some traffic sign must be 116 inches. Also, its length is 10 inches longer than its width. Find the dimensions of this sign. Please help me solve to obtain the width of the sign is ____ inches and the length of the sign is _____ inches.
----
Let width = x
Then length = x + 10
----
Perimeter = 2(length + width)
116 = 2(x+10 + x)
58 = 2x+10
2x = 48
x = 24 inches (width)
x+10 = 34 inches (length)
============================
Cheers,
Stan H.
==================